Kedia Inc. forecasts a negative free cash flow for the coming year, FCF1 = -$10 million, but it expects positive numbers thereafter, with FCF2 = $25 million. After Year 2, FCF is expected to grow at a constant rate of 4% forever. If the weighted average cost of capital is 14.0%, what is the firm's total corporate value, in millions?
Answer #1

FCF1=-$10

FCF2=$25

g=4%

WACC =14%

terminal value at t=2

HV2=FCF2(1+g)/(WACC-g)=25(1+0.04)/(0.14-0.04)

HV2=$260

Firms total value of operations =-10/(1+0.14)1+(25+260)/(1+0.14)2

Firms total value of operations =$210.53
Answer #2
Hi, Please find the answer as follows: TV = FCF3/(wacc - g) = 25*(1.04)/(.14-.04) = 260 =-10/ + 25/(1+.14)^2 + 260/(1+.14)^2 = 199.75 or 200 is the corporate value in millions. Thanks, Aman